# Energy Reflection and Transmission of Interfaces in $T\bar{T}$-deformed CFT

### v1

@@ 5 Energy transmission in the holographic NESS
This leaves us to determine $X_{L,R}(\sigma)$ from the remaining of the two equations, which then completely solves the gluing problem. However, these are ordinary differential equations and an exact analytic solution may be difficult to come up with. So we adapt an alternate strategy following Bachas:2021tnp. Firstly, we compute the determinant of the induced metric $h=-\text{det}h_{ab}$ from both sides and invert them to get
@@ 5.1 Extracting the transmission function
Now since $A>0$, it turns out from (67) that the worldsheet becomes spacelike as the strings enters into the the ergoregion $\sigma=0$, if $\sigma_{+}<0$. So for timelike worldsheet we must have $\sigma_{+}\geq 0$. The case with $\sigma_{+}>0$ corresponds to a turning point of the membrane Bachas:2021tnp and we typically do not want that since our discussion concerns about an isolated defect. For the other choice $\sigma_{+}=0$, the membrane enters into the ergoregion and remains timelike, and one can further show that it never comes out again Bachas:2021tnp. This condition amounts to $B>0$ and $C=0$. The expression for $C$ takes a much simpler form in terms of the variables
@@
where we are considering the folded setup of ICFT following Bachas:2021tnp, and we have used the fact that in the region $IIIA(B)$, the right movers are purely thermal at temperature $\theta_{L(R)}$. Using (71) back in (70) along with the relation $J_{R}=-J_{L}$, one can extract the transmission coefficient from the resulting Stefan-Boltzmann law $J_{L}=\frac{\pi c}{6}\mathcal{T}_{L}^{(0)}(\theta_{L}^{2}-\theta_{R}^{2})$ with
@@
6 Gravitational Scattering in deformed theory
In this section, we set up the gravitational scattering following Bachas:2020yxv to extract the reflection and transmission coefficients of a holographic interface gluing a pair of $T\bar{T}$-deformed CFTs. The scattering state is realized by shooting null waves from the asymptotic null infinity of the left (deformed) CFT and collecting the flux at the future null infinities of the two (deformed) CFTs, as described in section 2 (see Figure 1). In the holographic model, this boundary scattering can be realized by gluing two deformed spacetimes with boundary gravitons across a thin tensile string. Now as discussed in 3, these deformed geometries are obtained from the corresponding undeformed geometries with boundary graviton Bachas:2020yxv by inverting the map  [^9] Note that, the multiplicative factors in this map is slightly different compared to section 3. This is to relate to the results of Bachas:2020yxv.
@@ 6 Gravitational Scattering in deformed theory
where the integrands are related to the vevs of the boundary stress-tensor. However, the vevs in this case being coordinate-dependent, the inversion is non-trivial. So following Bachas:2020yxv (and also similar to section 4), we shall treat the vevs as infinitesimal excitation over the ICFT vacuum and invert the map perturbatively, leading to the deformed spacetimes
@@
which satisfies Einstein’s equation perturbatively. Note that the perturbative geometries know about the deformation even at $\mathcal{O}(\epsilon)$, although the boundary CFTs acquire a trace at $\mathcal{O}(\epsilon^{2})$. At linearized order, even in the deformed theory, the functions $\mathcal{L}$ and $\bar{\mathcal{L}}$ are proportional to the stress-tensor vevs (see Eq. (25)). Now following Bachas:2020yxv, we choose the following ansatz for the vevs:
@@ 6.2 First order
Next, while $\Delta(0)=0$ is trivially satisfied, $\zeta(0)=0$ and $\gamma(0)=0$ fix the coefficients $a_{+}(\omega)$ and $a_{-}(\omega)$ of the homogenous plane-wave solutions. Now, as discussed in Bachas:2020yxv, it turns out that the Israel gluing conditions along with the Dirichlet boundary conditions are not sufficient to extract the transmission coefficients and one needs additional conditions. In this context, the authors imposed a no-outgoing-wave condition at the Poincaré horizon, which in turn sets $a_{+}(\omega)=0$. In the present context, this leads to
@@
where in the second step have used (84)-(85). This precisely agrees with the undeformed transmission coefficient reported in Bachas:2020yxv. Next, we consider the zero-tension limit, which also requires $\theta_{L}=\theta_{R}=0$ from (84). In this limit, we have $\mathcal{T}^{(\mu)}_{L}=1$, which is expected as the interface becomes transparent when the tension vanishes. Finally, to compare with the flow equation results, we consider the equal central-charge limit, which amounts to setting $\theta_{R}=\theta_{L}$, and hence $\mu_{L}=\mu_{R}=\mu$ from (88). This gives
@@
As shown in Bachas:2020yxv, for an undeformed ICFT, the analysis to linear order in vev is sufficient to unambiguously fix the universal coefficient to all orders in the expansion. However, in the case of an interface in deformed CFT, (91) is expected to receive correction from the non-linear [^11] Note that, even for the deformed geometries, the Fefferman-Graham expansion truncates at second order, but the Israel gluing conditions do not. transmission coefficients resulting from the higher derivatives of the transmission function. To capture this, one needs to carry out the gluing beyond linear order and express the outgoing fluxes in terms of the incoming one using our modified matching condition (41). This would be technically more involved and we leave it for future work.
@@ 7 Conclusions and outlook
Let us summarize our work. In this paper, we have studied energy transport properties of an interface gluing a pair of $T\bar{T}$-deformed 1+1D holographic CFTs. Gluing of such CFTs must obey the relation $c_{L}\mu_{L}=c_{R}\mu_{R}$. Our approach has been two-fold. Firstly, we have used the flow equations of $T\bar{T}$ deformation to compute the vevs of the stress-tensor in the deformed theory. Since the flow equations are state-dependent, for the concreteness of our analysis, we chose two classes of states —  a NESS with a heat current and a scattering state. In this context, we have also proposed a non-linear modification of the usual gluing conditions, captured by a function of the incoming fluxes, which we call the transmission function. The non-linear matching conditions ensure continuity of current across the interface. For the equivalence of the two matching conditions under the exchange of incoming fluxes, the transmission function must be purely antisymmetric in its arguments.
@@
In a parallel approach, we have used the mixed-boundary condition interpretation of the deformation along with the thin-brane model of ICFT, to reproduce the flow equation results by performing a holographic computation in the bulk dual of these states. Towards that, we have given a rigorous account of gluing a pair of deformed geometries across the brane obeying the standard Israel junction conditions. For the NESS, we have reproduced the transmission function up to quartic order in incoming fluxes, whereas for the scattering state we have precisely reproduced the energy-dependent transmission coefficient from various holographic conditions. We can see these results as a partial generalization of the universality property of the energy transmission coefficient in the CFT, extended now to the non-linear (for the NESS) or frequency-dependent (for the scattering state) coefficients. As we have discussed, however, we cannot ascertain the universality in the sense of independence from the operator that creates the excitation.
@@
Several questions remain to be addressed. It would be interesting to have a closed-form expression for the transmission function for the NESS; this can perhaps be done by computing more orders in the perturbative expansion and attempting a resummation. It would be important to extend our analysis beyond linear order for the scattering state. The computation becomes more involved, and it is likely that the condition of current continuity has to be generalized to account for the non-vanishing trace of the energy-momentum tensor (as was shown in the case of a relevant deformation of a topological defect in Bajnok:2013waa). It would also be interesting to see if our results could be reproduced by the methods of generalized hydrodynamics Castro-Alvaredo:2016cdj; Piroli:2017ltm, imposing a condition at the interface for the scattering of quasi-particles. The most important question is whether our results can be extended beyond the holographic context to the full quantum theory. Finally, it would be interesting to repeat the holographic computation using conformal boundary condition at the cutoff surface; this prescription has been argued to correspond to coupling the CFT to a timelike Liouville theory and deforming by a Liouville-dressed $T\bar{T}$ operator Allameh:2025gsa; the resulting theory is expected to have a better UV behavior, and it would be interesting to see if this would cure the violation of unitarity we found in the high-frequency regime of the transmission coefficient. [^12] We thank Dionysios Anninos for the suggestion.

### v2

@@ 5 Energy transmission in the holographic NESS
This leaves us to determine $X_{L,R}(\sigma)$ from the remaining of the two equations, which then completely solves the gluing problem. However, these are ordinary differential equations and an exact analytic solution may be difficult to come up with. So we adapt an alternate strategy following Bachas et al. (2021). Firstly, we compute the determinant of the induced metric $h=-\text{det}h_{ab}$ from both sides and invert them to get
@@ 5.1 Extracting the transmission function
Now since $A>0$, it turns out from (67) that the worldsheet becomes spacelike as the strings enters into the the ergoregion $\sigma=0$, if $\sigma_{+}<0$. So for timelike worldsheet we must have $\sigma_{+}\geq 0$. The case with $\sigma_{+}>0$ corresponds to a turning point of the membrane Bachas et al. (2021) and we typically do not want that since our discussion concerns about an isolated defect. For the other choice $\sigma_{+}=0$, the membrane enters into the ergoregion and remains timelike, and one can further show that it never comes out again Bachas et al. (2021). This condition amounts to $B>0$ and $C=0$. The expression for $C$ takes a much simpler form in terms of the variables
@@
where we are considering the folded setup of ICFT following Bachas et al. (2021), and we have used the fact that in the region $IIIA(B)$, the right movers are purely thermal at temperature $\theta_{L(R)}$. Using (71) back in (70) along with the relation $J_{R}=-J_{L}$, one can extract the transmission coefficient from the resulting Stefan-Boltzmann law $J_{L}=\frac{\pi c}{6}\mathcal{T}_{L}^{(0)}(\theta_{L}^{2}-\theta_{R}^{2})$ with
@@
6 Gravitational Scattering in the deformed spacetime
In this section, we set up the gravitational scattering following Bachas et al. (2020) to extract the reflection and transmission coefficients of a holographic interface gluing a pair of $T\bar{T}$-deformed CFTs. The scattering state is realized by shooting null waves from the asymptotic null infinity of the left (deformed) CFT and collecting the flux at the future null infinities of the two (deformed) CFTs, as described in section 2 (see Figure 1). In the holographic model, this boundary scattering can be realized by gluing two deformed spacetimes with boundary gravitons across a thin tensile string. Now as discussed in 3, these deformed geometries are obtained from the corresponding undeformed geometries with boundary graviton Bachas et al. (2020) by inverting the map  [^8] Note that, the multiplicative factors in this map is slightly different compared to section 3. This is to relate to the results of Bachas et al. (2020).
@@ 6 Gravitational Scattering in the deformed spacetime
where the integrands are related to the vevs of the boundary stress-tensor. However, the vevs in this case being coordinate-dependent, the inversion is non-trivial. So following Bachas et al. (2020) (and also similar to section 4), we shall treat the vevs as infinitesimal excitation over the ICFT vacuum and invert the map perturbatively, leading to the deformed spacetimes
@@
which satisfies Einstein’s equation perturbatively. Note that the perturbative geometries know about the deformation even at $\mathcal{O}(\epsilon)$, although the boundary CFTs acquire a trace at $\mathcal{O}(\epsilon^{2})$. At linearized order, even in the deformed theory, the functions $\mathcal{L}$ and $\bar{\mathcal{L}}$ are proportional to the stress-tensor vevs (see Eq. (25)). Now following Bachas et al. (2020), we choose the following ansatz for the vevs:
@@ 6.2 First order
Next, while $\Delta(0)=0$ is trivially satisfied, $\zeta(0)=0$ and $\gamma(0)=0$ fix the coefficients $a_{+}(\omega)$ and $a_{-}(\omega)$ of the homogenous plane-wave solutions. Now, as discussed in Bachas et al. (2020), it turns out that the Israel gluing conditions along with the Dirichlet boundary conditions are not sufficient to extract the transmission coefficients and one needs additional conditions. In this context, the authors imposed a no-outgoing-wave condition at the Poincaré horizon, which in turn sets $a_{+}(\omega)=0$. In the present context, this leads to
@@
where in the second step have used (84)-(85). This precisely agrees with the undeformed transmission coefficient reported in Bachas et al. (2020). Next, we consider the zero-tension limit, which also requires $\theta_{L}=\theta_{R}=0$ from (84). In this limit, we have $\mathcal{T}^{(\mu)}_{L}=1$, which is expected as the interface becomes transparent when the tension vanishes. Finally, to compare with the flow equation results, we consider the equal central-charge limit, which amounts to setting $\theta_{R}=\theta_{L}$, and hence $\mu_{L}=\mu_{R}=\mu$ from (88). This gives
@@
As shown in Bachas et al. (2020), for an undeformed ICFT, the analysis to linear order in vev is sufficient to unambiguously fix the universal coefficient to all orders in the expansion. However, in the case of an interface in deformed CFT, (91) is expected to receive correction from the non-linear [^10] Note that, even for the deformed geometries, the Fefferman-Graham expansion truncates at second order, but the Israel gluing conditions do not. transmission coefficients resulting from the higher derivatives of the transmission function. To capture this, one needs to carry out the gluing beyond linear order and express the outgoing fluxes in terms of the incoming one using our modified matching condition (41). This would be technically more involved and we leave it for future work.
@@ 7 Conclusions and outlook
Let us summarize our work. In this paper, we have studied energy transport properties of an interface gluing a pair of $T\bar{T}$-deformed 1+1D holographic CFTs. Gluing of such CFTs must obey the relation $c_{L}\mu_{L}=c_{R}\mu_{R}$. Our approach has been two-fold. Firstly, we have used the flow equations of $T\bar{T}$ deformation to compute the vevs of the stress-tensor in the deformed theory. Since the flow equations are state-dependent, for the concreteness of our analysis, we chose two classes of states —  a NESS with a heat current and a scattering state. In this context, we have also proposed a non-linear modification of the usual gluing conditions, captured by a function of the incoming fluxes, which we call the transmission function. The non-linear matching conditions ensure continuity of current across the interface. Furthermore, the self-consistency of the matching conditions under the exchange of the incoming fluxes requires the transmission function to be purely antisymmetric in its arguments. For the scattering state, the non-trivial temporal dependence of the fluxes restricts our analysis to linear order only, wherein the transmission function can be equivalently recast as a frequency-dependent transmission coefficient. This in turn puts an upper bound on the allowed frequencies following from the unitarity constraints.
@@
In a parallel approach, we have used the mixed-boundary condition interpretation of the deformation along with the thin-brane model of ICFT, to reproduce the flow equation results by performing a holographic computation in the bulk dual of these states. Towards that, we have given a rigorous account of gluing a pair of deformed geometries across the brane obeying the standard Israel junction conditions. For the NESS, we have reproduced the transmission function up to quartic order in incoming fluxes, whereas for the scattering state we have precisely reproduced the energy-dependent transmission coefficient from various holographic conditions. In We can see these results as a partial generalization of the universality property of the energy transmission coefficient in the CFT, extended now to the non-linear (for the NESS) or frequency-dependent (for the scattering state) coefficients. As we have discussed, however, we cannot ascertain the universality in the sense of independence from the operator that creates the excitation.
@@
Several questions remain to be addressed. It would be interesting to have a closed-form expression for the transmission function for the NESS; this can perhaps be done by computing more orders in the perturbative expansion and attempting a resummation. It would be important to extend our analysis beyond linear order for the scattering state. The computation becomes more involved, and it is likely that the condition of current continuity has to be generalized to account for the non-vanishing trace of the energy-momentum tensor (as was shown in the case of a relevant deformation of a topological defect in Bajnok et al. (2014)). It would also be interesting to see if our results could be reproduced by the methods of generalized hydrodynamics Castro-Alvaredo et al. (2016); Piroli et al. (2017), imposing a condition at the interface for the scattering of quasi-particles. The most important question is whether our results can be extended beyond the holographic context to the full quantum theory. Also, it would be worth exploring how does the (modified) energy transmsission properties of the interface affect the density of states of the two deformed CFTs. In particular, it would be interesting to investigate the existence of negative specific heat in the deformed CFTs Barbon and Rabinovici (2020) in the presence of the interface. This perhaps can most neatly be addressed by putting the theory on a finite volume, which however, would require a pair of interfaces rather than one. Finally, it would be interesting to repeat the holographic computation using conformal boundary condition at the cutoff surface; this prescription has been argued to correspond to coupling the CFT to a timelike Liouville theory and deforming by a Liouville-dressed $T\bar{T}$ operator Allameh and Shaghoulian (2025); the resulting theory is expected to have a better UV behavior, and it would be interesting to see if this would cure the violation of unitarity we found in the high-frequency regime of the transmission coefficient. [^11] We thank Dionysios Anninos for the suggestion.